OpenStudy (anonymous):

Kinetic energy KE=1/2mv^2 if mass is increased by 4% And velocity reduced by 3% determine the approximate change in K.E... Apparently i need to use binomial theorem (a=x)^n but how do i do it?

OpenStudy (anonymous):

Its really simple. New mass = 1+4/100 i.e\[Mass_n = 1 + \frac{ 4 }{ 100 }\]

OpenStudy (anonymous):

For velocity subtract \[Velocity_n = 1-0.3 = 0.97\]

OpenStudy (anonymous):

Square .97 now

OpenStudy (anonymous):

Tell me you get .

OpenStudy (anonymous):

For distinction I need to use binomial.... Thank you. I think i have done something like that.

OpenStudy (anonymous):

The Binomial theorem example: (x+y)^2 = x^2 + y^2 + 2xy

OpenStudy (anonymous):

How can we possible apply the Binomial theorem here..??

OpenStudy (anonymous):

I don't know. Will have to check with my teacher... Thank you for your help.

OpenStudy (michele_laino):

If I call with \(v_1,m_1\) the new values of the speed and mass, and with \(v_0,m_0\) the corresponding initial value, then I can write this: \[\large \begin{gathered} {m_1} = 1.04{m_0} \hfill \\ {v_1} = 0.97{v_0} \hfill \\ \hfill \\ \frac{{\frac{1}{2}{m_0}v_0^2 - \frac{1}{2}{m_1}v_1^2}}{{\frac{1}{2}{m_0}v_0^2}} = \frac{{{m_0}v_0^2 - 1.04{m_0} \cdot {{\left( {0.97} \right)}^2}v_0^2}}{{{m_0}v_0^2}} = \hfill \\ \hfill \\ = \frac{{{m_0}v_0^2\left\{ {1 - 1.01 \cdot {{\left( {0.97} \right)}^2}} \right\}}}{{{m_0}v_0^2}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

values*

OpenStudy (michele_laino):

oops.. I have made a typo: \[\large \begin{gathered} \frac{{\frac{1}{2}{m_0}v_0^2 - \frac{1}{2}{m_1}v_1^2}}{{\frac{1}{2}{m_0}v_0^2}} = \frac{{{m_0}v_0^2 - 1.04{m_0} \cdot {{\left( {0.97} \right)}^2}v_0^2}}{{{m_0}v_0^2}} = \hfill \\ \hfill \\ = \frac{{{m_0}v_0^2\left\{ {1 - 1.04 \cdot {{\left( {0.97} \right)}^2}} \right\}}}{{{m_0}v_0^2}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (alekos):

i see what your doing this is like finding ΔΕ/Ε0 correct ?

OpenStudy (anonymous):

Thank you. .

OpenStudy (michele_laino):

that's right!! @alekos